23. How many photons are emitted per second by a 5 mW laser source operating at 632.8 nm? 

Grade:12

1 Answers

Arun
6 years ago
the energy of each photon is = 
 E = hc / \lambda
 = 6.63 * 10 –34 * 3 * 108 / 632.8 * 10 –9
 = 3.14 * 10 –19  J
the energy emitted per second by the laser is =  5 * 10 –3  J
hence totla no. of photons emitted per second  = 5 * 10 –3  J / 3.14 * 10 –19  J
 = 1.6 * 1016